OpenStudy (alfordabby):

The equation of line ab is y=3x-2. write an equation of a line parallel to line ab in slope-intercept form that contains points (-2,4)

OpenStudy (alfordabby):

will fan & medal :)

OpenStudy (anonymous):

Parallel lines have the same slope

OpenStudy (anonymous):

(y-y1)=m(x-x1)

OpenStudy (alfordabby):

thank you. :)

OpenStudy (alfordabby):

is that the answer?

OpenStudy (anonymous):

No that is used to solve it...

OpenStudy (alfordabby):

is (3-3^1)m=(-2+3^1) the next step?

OpenStudy (alfordabby):

=m**

OpenStudy (anonymous):

4= y1 and -2 =x1 Your slope (m) is the same as the equation it is parallel to which is 3....So plug into the equation (y-y1)=m(x-x1) and then solve in slope intercept form